OpenStudy (anonymous):

the sum of two numbers that differ by 3 is x. In terms of x, what is the value of the greater of the two numbers? A)x-3/2 B)x/2 C)x+3/x D)x/2+3 E.2x-3/2

OpenStudy (accessdenied):

I would call the first number \(a - 3\), and subsequently the second number would be \(a\), so that \((a) - (a - 3) = 3\) We would then be able to create the equation: \((a - 3) + a = x\), or more simply, \(2a - 3 = x\), based on the given information. Then, we just solve for the larger number of the two, \(a\), leaving it in terms of \(x\). \[\quad 2a - 3 = x \\ \quad 2a = x + 3 \\ \quad a = \frac{x + 3}{2}\] Which is apparently not an answer choice? '\('o' )/`